For what value(s) of x ist the expression (x^2-4x+4)/(2x^2-3x-2) undifined?
Thanks

Answer by venugopalramana(3286)

For what value(s) of x ist the expression (x^2-4x+4)/(2x^2-3x-2) undifined?
WHEN DR=0
2X^2-4X+X-2=0
2X(X-2)+1(X-2)=0
(X-2)(2X+1)=0
X-2=0.....THAT IS ....X=2....AND ...2X+3=0....THAT IS .....X=-3/2